# Interference of photons from independent hot atoms

### v2

@@ Abstract
The coherence of light from independent ensembles of elementary atomic emitters plays a paramount role in diverse areas of modern optics. We demonstrate the interference of photons scattered from independent ensembles of warm atoms in atomic vapor. relies on the feasibility of the preservation of coherence of light scattered elastically in the forward and backward directions from Doppler-broadened atomic ensembles, such that photons with chaotic photon statistics from two opposite atomic velocity groups contribute to the same detection mode. While the random phase fluctuations of the scattered light caused by a large thermal motion prevent direct observability of the interference in the detected photon rate, the stable frequency difference between photons collected from scattering off counter-propagating laser beams provides strong periodic modulation of the photon coincidence rate with the period given by the detuning of the excitation laser from the atomic resonance. Presented interferometry promises direct applications in Doppler-free atomic and molecular spectroscopy.
@@ I Introduction
Optical coherence is paramount to all areas of modern optics. The feasibility of the coherence of light scattered from atoms enabled a broad range of essential applications ranging from optical atomic and molecular spectroscopy to diverse atomic sensors, efficient light-matter interfaces, and long-distance quantum communication [1, 2, 3, 4, 5, 6]. The corresponding experimental implementations and their theoretical analysis consistently acknowledge the detrimental role of thermal atomic motion, often corresponding to a predominant source of decoherence  [7, 8, 9, 10, 11, 12, 13, 14]. The usual approaches towards mitigation of its impact correspond to suppression of thermal kinetic energy by means of laser cooling, the employment of tight atomic trapping potentials, optimized excitation geometries, or diverse spatial and spectral filtering schemes for minimization of the which-way information due to atomic recoils.
@@
We demonstrate the coherence of light scattering in hot atomic vapors within a broad range of detunings from atomic resonance with a positive contribution of different atomic velocity classes. The coherence can be maintained both in the forward and backward scattering, where the latter provides a stable frequency difference between the scattered fields and the employed atomic transition. The verification of coherence using the interferometric scheme including light fields scattered from counter-propagating excitation laser beams provides intrinsic applications in atomic spectroscopy. It allows for a direct estimation of the absolute value of the laser detuning from atomic resonance without any need for the modulation of the laser probe frequency or atomic transition parameters, and neither requires calibration of any interaction parameters or thermal atomic motion.
II Interference scheme
@@
$E(t)\propto\sum_{i=1}^{N}E_{i}e^{-\mathrm{i}(\omega_{\mathrm{F},i}t-\textbf{k}_{\mathrm{F}}\cdot\textbf{r}_{i})}+\sum_{j=1}^{N}E_{j}e^{-\mathrm{i}(\omega_{\mathrm{B},j}t-\textbf{k}_{\mathrm{B}}\cdot\textbf{r}_{j})}.$ (1)
where $E_{i\,(j)}$ are the amplitudes of fields scattered from $i\,(j)$-th atoms, $\mathbf{r}$ are the position vectors of atoms. The summation includes contributions from a large number of $N\gg 1$ atoms in each of the two opposite atomic velocity classes. Here, we assume that the numbers of contributing atoms to the forward and backward scattered fields are approximately the same with relevance to the presented experimental configuration, where the numbers of contributing atoms are given by the observation volume of the two atomic ensembles with the same atomic density. The Doppler distribution of atomic velocity classes of thermal vapor affects the number of contributing atoms in both ensembles equally, as the effective laser detuning from atomic resonance is the same for both scattering contributions. The evaluation of the normalized second-order correlation function in the limit of thermally fluctuating atomic positions gives
@@
Here, the first-order temporal correlation functions $g^{(1)}_{\mathrm{F,B}}(\tau)$ reflect the coherence of light scattered on the individual atomic ensembles. The term $\bar{n}_{\mathrm{F(B)}}$ correspond to mean photon numbers from a particular atomic ensemble, which can be affected by different atomic level populations and losses experienced by the two fields. The complete derivation can be found in Supplementary Materials. The equation reveals the crucial contribution of beating of the first-order correlation functions $g^{(1)}(\tau)$, which modulates the second-order photon correlations at their mean frequency difference $\omega_{\mathrm{F}}-\omega_{\mathrm{B}}$. It represents an extended version of the Siegert relation for chaotic light with two contributing atomic thermal light sources at different frequencies [33]. The visibility of the beating of first-order coherences can be controlled by the relative contributions of two atomic velocity classes, which depends on the amount of detected forward and backward scattered light. For equal number of photons from the two atomic velocity classes contributing to the signal, the contrast is maximal. We emphasize that the resonant nature of the interaction is essential for the velocity selectivity of scatterers, as assumed in Eq. (1). Consequently, high visibility of interference requires $(\omega_{\mathrm{F}}-\omega_{\mathrm{B}})\approx 2\Delta>\sigma_{\mathrm{F}},\sigma_{\mathrm{B}}$, such that the emitted photons correspond to different non-overlapping spectral components with width given by the residual Doppler-broadening.
@@ III Experimental demonstration
Examples of the measured interference of light from independent warm atomic ensembles are shown in Fig. 3. The independent photon correlation measurements of forward and backward scattered fields for laser detuning $\Delta_{\mathrm{L}}\approx 100$ MHz in a) demonstrate a close-to ideal bunching values $g^{(2)}_{\mathrm{F}}(0)=1.94\pm 0.02$ and $g^{(2)}_{\mathrm{B}}(0)=1.92\pm 0.03$, signifying their chaotic photon statistics [33, 34, 32]. Importantly, these individual correlation functions demonstrate large temporal coherences corresponding to spectral bandwidths $\sigma_{\mathrm{F}}\approx 9.6$ MHz and $\sigma_{\mathrm{B}}\approx 15.8$ MHz, respectively. These spectral widths set an upper limit on the observable period of interference modulation in the corresponding $g^{(2)}_{\mathrm{F+B}}(\tau)$. They agree with their independent evaluations within the uncertainty of measurement of the scattering angle $\theta$ and confirm the predicted scaling of temporal coherence for the backward-scattered signal. We note that the presented backscattering measurements correspond to a first experimental study of $g^{(2)}_{\mathrm{B}}(\tau)$ from warm atoms in such configuration and provide an important confirmation of the feasibility of coherence of light in this direction. The plot in b) depicts the corresponding $g^{(2)}_{\mathrm{F+B}}(\tau)$, where both forward and backward-scattered fields contribute simultaneously. The interference results in clear modulation of photon coincidence detection probabilities for $|\tau|>0$ with the period of $f_{\mathrm{mod}}=210.8\pm 1.2$ MHz, in agreement with the independently set laser detuning from the $F=2\leftrightarrow F^{\prime}=3$. The periodic modulation between constructive and destructive interference at the frequency difference of interfering photons signifies the fulfillment of the critical conditions for the presented scheme. It confirms the sufficiency of the velocity selectivity in the given excitation regime and, at the same time, the mutually coherent contribution with sufficient coherence time from atoms at different velocity classes in a strongly Doppler-broadened thermal atomic ensemble. The detected mean photon rate for presented excitation conditions was $R_{\mathrm{(F+B)}}=(64\pm 2)\times 10^{4}$ counts/s. The measured photon bunching at zero time delay $g^{(2)}_{\mathrm{F+B}}(0)=1.96\pm 0.01$ and the rate of simultaneous detection of two photons $R_{\mathrm{C}}=(4.9\pm 0.3)\times 10^{3}$ coincidences/s for the time bin of 14.3 ns.
@@ IV Conclusions
The presented experimental methodology provides an exclusive tool for observation of the emergence of coherence of light from independent ensembles of elementary atomic scatterers. It is directly applicable in a broad range of fundamental interferometric tests and in applications in optical atomic and molecular spectroscopy [1, 36, 37, 38, 39]. It can be implemented to different classical or quantum scatterers with feasible resonant interaction, independently of the availability of their motional cooling. Importantly, contrary to the broad range of spectroscopy methods based on a free-space laser absorption with sub-Doppler resolution applicable to warm atoms, the presented interference scheme allows for implementation with small samples of atomic or molecular vapors and with electronic transitions providing only very low resonant optical depth. Despite operating in a small saturation parameter limit, the resonant character of interaction is indispensable for its functioning, as it provides the necessary selectivity of velocity classes of contributing scatterers. For atoms, the observation of interference can thus be considered as a unique signature of the quantized internal structure in the elastic scattering limit. Importantly, the interference is not compromised by the large thermal motion. On the contrary, it enables a large measurement bandwidth of atomic spectroscopy. The combination of a high atomic velocity selectivity and photon correlation detection provides sub-Doppler and sub-natural linewidth precision in the regime of large thermal motion. The high atomic temperature enhances the spectral bandwidth of the spectroscopy and the temporal resolution and stability of frequency estimation can benefit from the availability of a large number of atoms in vapor cells. Extensions of presented elementary examples to atoms or molecules with closely spaced electronic level manifolds can result in complex beating structures providing information about the contributing energy level structure.
Acknowledgements.
@@
S. J. and J. M. acknowledge the support of the Czech Science Foundation under the project GA21-13265X. L. S. is grateful for the national funding from the MEYS under the project CZ.02.01.01/00/22 008/0004649. L. S. and R. K. acknowledge funding from the QUANTERA ERA-NET cofund in quantum technologies implemented within the European Union’s Horizon 2020 Programme (project PACE-IN, 8C20004). L. S. is grateful to Morgan Mitchell for stimulating discussion on the scattering of light in warm atomic vapors.
@@
$E(t)\propto\sum_{i=1}^{N}E_{i}e^{-\mathrm{i}(\omega_{{\rm F},i}t-\textbf{k}_{\rm F}\cdot\textbf{r}_{i})}+\sum_{j=1}^{N}E_{j}e^{-\mathrm{i}(\omega_{{\rm B},j}t-\textbf{k}_{\rm B}\cdot\textbf{r}_{j})}$ (S1)
where $E_{i,(j)}$ is the amplitude of the fields from $i(j)$-th atom, $\mathbf{k}$ are the wave vectors of scattered light, and $\mathbf{r}$ are the atomic position vectors. The summation includes contributions from $N$ atoms in each of the two velocity classes. With relevance to the presented experimental configuration, the number of contributing atoms to the forward and backward scattered fields is assumed to be equal. It is determined by the excitation-observation mode overlap of the two atomic ensembles, which is set mostly by the same diameter of the excitation laser beam and of the observation Gaussian mode. In addition, the Doppler distribution of atomic velocity classes of thermal vapor affects the number of contributing atoms in both ensembles equally, as the effective laser detuning is the same for both forward and backward scattered contributions. The evaluation of the second-order correlation function

### v3

@@ Abstract
The coherence of light from independent ensembles of elementary atomic emitters plays a paramount role in diverse areas of modern optics. We demonstrate the interference of photons scattered from independent ensembles of warm atoms in atomic vapor. It relies on the feasibility of the preservation of coherence of light scattered elastically in the forward and backward directions from Doppler-broadened atomic ensembles, such that photons with chaotic photon statistics from two opposite atomic velocity groups contribute to the same detection mode. While the random phase fluctuations of the scattered light caused by a large thermal motion prevent direct observability of the interference in the detected photon rate, the stable frequency difference between photons collected from scattering off counter-propagating laser beams provides strong periodic modulation of the photon coincidence rate with the period given by the detuning of the excitation laser from the atomic resonance. The presented interferometry represents a sensitive and robust methodology for Doppler-free optical atomic and molecular spectroscopy based on photon correlation measurements on scattered light.
@@ I Introduction
Optical coherence is paramount to all areas of modern optics. The feasibility of the coherence of light scattered from atoms enabled a broad range of essential applications ranging from optical atomic and molecular spectroscopy to diverse atomic sensors, efficient light-matter interfaces, and long-distance quantum communication [1, 2, 3, 4, 5, 6]. The corresponding experimental approaches and their theoretical analysis consistently acknowledge severely detrimental impact of thermal atomic motion as a predominant source of decoherence [7, 8, 9, 10, 11, 12, 13, 14]. The usual approaches towards mitigation of its impact correspond to suppression of thermal kinetic energy by means of laser cooling, the employment of tight atomic trapping potentials, optimized excitation geometries, or diverse spatial and spectral filtering schemes for minimization of the which-way information due to atomic recoils.
@@
We demonstrate the coherence of light scattering in hot atomic vapors within a broad range of detunings from atomic resonance with a positive contribution of different atomic velocity classes. The coherence can be maintained both in the forward and backward scattering, where the latter provides a stable frequency difference between the scattered fields and the employed atomic transition. The verification of coherence using the interferometric scheme including light fields scattered from counter-propagating excitation laser beams reveals intrinsic application in atomic spectroscopy. The methodology is based on the measurement of atomic fluorescence in photon correlation signal and allows for a sub-Doppler spectroscopy in low optical depth regime applicable to dilute ensembles of hot atoms. It allows for a direct estimation of the absolute value of the laser detuning from atomic resonance without any need for the modulation of the laser probe frequency or atomic transition parameters, and neither requires calibration of any interaction parameters or thermal atomic motion.
II Interferometry with hot atomic scatterers
@@
$E(t)\propto\sum_{i=1}^{N}E_{{\rm F},i}e^{-\mathrm{i}(\omega_{\mathrm{F},i}t-\textbf{k}_{\mathrm{F}}\cdot\textbf{r}_{i})}+\sum_{j=1}^{N}E_{{\rm B},j}e^{-\mathrm{i}(\omega_{\mathrm{B},j}t-\textbf{k}_{\mathrm{B}}\cdot\textbf{r}_{j})}.$ (1)
where $E_{{\rm F},i({\rm B},j)}$ are the amplitudes of the fields from $i(j)$-th atom, $\mathbf{r}$ are the position vectors of atoms. The summation includes contributions from a large number of $N\gg 1$ atoms in each of the two opposite atomic velocity classes. Here, we assume that the numbers of contributing atoms to the forward and backward scattered fields are approximately the same with relevance to the presented experimental configuration, where the numbers of contributing atoms are given by the observation volume of the two atomic ensembles with the same atomic density. The Doppler distribution of atomic velocity classes of thermal vapor affects the number of contributing atoms in both ensembles equally, as the effective laser detuning from atomic resonance is the same for both scattering contributions. The evaluation of the normalized second-order correlation function in the limit of thermally fluctuating atomic positions gives
@@
Here, the first-order temporal correlation functions $g^{(1)}_{\mathrm{F,B}}(\tau)$ reflect the coherence of light scattered on the individual atomic ensembles. The term $\bar{n}_{\mathrm{F(B)}}$ correspond to mean photon number from a particular atomic ensemble, which can be affected by different atomic level populations and losses experienced by the two fields. The complete derivation can be found in Supplementary Materials. The equation reveals the crucial contribution of beating of the first-order correlation functions $g^{(1)}(\tau)$, which modulates the second-order photon correlations at their mean frequency difference $\omega_{\mathrm{F}}-\omega_{\mathrm{B}}$. It represents an extended version of the Siegert relation for chaotic light with two contributing atomic thermal light sources at different frequencies [33]. The visibility of the beating of first-order coherences can be controlled by the relative contributions of two atomic velocity classes, which depends on the amount of detected forward and backward scattered light. For equal number of photons from the two atomic velocity classes contributing to the signal, the contrast is maximal. We emphasize that the resonant nature of the interaction is essential for the velocity selectivity of scatterers, as assumed in Eq. (1). Consequently, high visibility of interference requires $(\omega_{\mathrm{F}}-\omega_{\mathrm{B}})\approx 2\Delta>\sigma_{\mathrm{F}},\sigma_{\mathrm{B}}$, such that the emitted photons correspond to different non-overlapping spectral components with width given by the residual Doppler-broadening.
@@ III Experimental demonstration
Examples of the measured interference of light from independent warm atomic ensembles are shown in Fig. 3. The independent photon correlation measurements of forward and backward scattered fields for laser detuning $\Delta_{\mathrm{L}}\approx 100$ MHz in a) demonstrate a close-to ideal bunching values $g^{(2)}_{\mathrm{F}}(0)=1.94\pm 0.02$ and $g^{(2)}_{\mathrm{B}}(0)=1.92\pm 0.03$, emerging from the random phase shifts of scattered photons caused by the large thermal motion of atoms [33, 34, 32]. Importantly, these individual correlation functions demonstrate large temporal coherences corresponding to spectral bandwidths $\sigma_{\mathrm{F}}\approx 9.6$ MHz and $\sigma_{\mathrm{B}}\approx 15.8$ MHz, respectively. They effectively correspond to average rates of random phase changes in each of the atomic ensembles and set an upper limit on the observable period of interference modulation in the corresponding $g^{(2)}_{\mathrm{F+B}}(\tau)$. The spectral widths agree with their independent evaluations within the uncertainty of measurement of the scattering angle $\theta$ and confirm the predicted scaling of temporal coherence for the backward-scattered signal. We note that the presented backscattering measurements correspond to a first experimental study of $g^{(2)}_{\mathrm{B}}(\tau)$ from warm atoms in such configuration and provide an important confirmation of the feasibility of coherence of light in this direction. The plot in b) depicts the corresponding $g^{(2)}_{\mathrm{F+B}}(\tau)$, where both forward and backward-scattered fields contribute simultaneously. The interference results in clear modulation of photon coincidence detection probabilities for $|\tau|>0$ with the period of $f_{\mathrm{mod}}=210.8\pm 1.2$ MHz, in agreement with the independently set laser detuning from the $F=2\leftrightarrow F^{\prime}=3$. The periodic modulation between constructive and destructive interference at the frequency difference of interfering photons signifies the fulfillment of the critical conditions for the presented scheme. It confirms the sufficiency of the velocity selectivity in the given excitation regime and, at the same time, the mutually coherent contribution with sufficient coherence time from atoms at different velocity classes in a strongly Doppler-broadened thermal atomic ensemble. The detected mean photon rate for presented excitation conditions was $R_{\mathrm{(F+B)}}=(64\pm 2)\times 10^{4}$ counts/s. The measured photon bunching at zero time delay $g^{(2)}_{\mathrm{F+B}}(0)=1.96\pm 0.01$. It allows for evaluating an upper limit on the ratio $r=0.02\pm 0.005$ between the mean photon number of the residual phase-correlated light and the light from the two independent chaotic light sources with mutually randomized phases. The derivation of the simple model for quantifying residual phase correlations can be found in the Supplementary Information. The rate of simultaneous detection of two photons $R_{\mathrm{C}}=(4.9\pm 0.3)\times 10^{3}$ coincidences/s for the time bin of 14.3 ns.
@@ IV Conclusions
The presented experimental methodology provides an exclusive tool for observation of the emergence of coherence of light from independent ensembles of elementary atomic scatterers. It is directly applicable in a broad range of fundamental interferometric tests and in applications in optical atomic and molecular spectroscopy [1, 36, 37, 38, 39]. It can be implemented to different classical or quantum scatterers with feasible resonant interaction, independently of the availability of their motional cooling. Importantly, contrary to the broad range of spectroscopy methods based on a free-space laser absorption with sub-Doppler resolution applicable to warm atoms, the presented interference scheme allows for implementation with small samples of atomic or molecular vapors and with electronic transitions providing only very low resonant optical depth. Despite operating in a small saturation parameter limit, the resonant character of interaction is indispensable for its functioning, as it provides the necessary selectivity of velocity classes of contributing scatterers. For atoms, the observation of interference can thus be considered as a unique signature of the quantized internal structure in the elastic scattering limit. Importantly, the interference is not compromised by the large thermal motion. On the contrary, it enables a large measurement bandwidth of atomic spectroscopy. The combination of a high atomic velocity selectivity and photon correlation detection provides sub-Doppler and sub-natural linewidth precision in the regime of large thermal motion with demonstrated feasibility of implementation at the single photon level. The high atomic temperature enhances the spectral bandwidth of the spectroscopy, which differentiates from other laser spectroscopy methods applicable to hot atoms [1]. While the temporal resolution and stability of frequency estimation could in the presented example benefit from the large number of Rb atoms in a vapor cell, the scheme can be particularly useful for analysis of extremely small and dilute atomic or molecular samples due to its intrinsic stability stemming from the observation of the interference signal in second-order correlations. In contrast to the schemes employing direct observation of the first-order interference, the presented methodology is not affected by slow phase drifts of the excitation laser and of interfering signals. These features promise direct applicability for precision optical spectroscopy of dipole transitions of atomic vapors including only about $10^{5}$ atoms in the observation mode, which has the potential to enhance laser spectroscopy of diverse rare atomic isotopes with applicable energy level structure, for example, 46Ca, with relevance to applications [40, 41]. Extensions of presented elementary examples to atoms or molecules with closely spaced electronic level manifolds can result in complex beating structures providing information about the contributing energy level structure.
V Acknowledgments
@@
S. J. and J. M. acknowledge the support of the Czech Science Foundation under the project GA21-13265X. L. S. is grateful for the national funding from the MEYS under the project CZ.02.01.01/00/22 008/0004649. L. S. and R. K. acknowledge funding from the QUANTERA ERA-NET cofund in quantum technologies implemented within the European Union’s Horizon 2020 Programme (project PACE-IN, 8C20004). L. S. is grateful to Romain Bachelard for stimulating discussions.
@@
$E(t)\propto\sum_{i=1}^{N}E_{{\rm F},i}e^{-\mathrm{i}(\omega_{{\rm F},i}t-\textbf{k}_{\rm F}\cdot\textbf{r}_{i})}+\sum_{j=1}^{N}E_{{\rm B},j}e^{-\mathrm{i}(\omega_{{\rm B},j}t-\textbf{k}_{\rm B}\cdot\textbf{r}_{j})}$ (S1)
where $E_{{\rm F},i({\rm B},j)}$ are the amplitudes of the fields from $i(j)$-th atom, $\mathbf{k}$ are the wave vectors of scattered light, and $\mathbf{r}$ are the atomic position vectors. The summation includes contributions from $N$ atoms in each of the two velocity classes. With relevance to the presented experimental configuration, the number of contributing atoms to the forward and backward scattered fields is assumed to be equal. It is determined by the excitation-observation mode overlap of the two atomic ensembles, which is set mostly by the same diameter of the excitation laser beam and of the observation Gaussian mode. In addition, the Doppler distribution of atomic velocity classes of thermal vapor affects the number of contributing atoms in both ensembles equally, as the effective laser detuning is the same for both forward and backward scattered contributions. The evaluation of the second-order correlation function
@@
S0.3 Model of radiation with partial phase correlations
We introduce a simple model describing the reduction of the second-order correlation function $g^{(2)}(0)$ by considering mixing of the light scattered from warm atoms with the phase-correlated light corresponding to the coherent state $|\alpha\rangle$. It allows for evaluation of a lower limit of the contribution of light from scatterers with mutually fully randomized phases of their optical field contributions to the detected mode. This phase randomization of excitation lasers secures the independence of the two interfering fields. As described in the main manuscript part, it emerges from different contributions of thermal atomic motion, manifesting in random atomic positions and Doppler shifts, which result in effective spectral widths $\sigma_{\mathrm{F}},\sigma_{\mathrm{B}}$ of the interfering light fields. The introduction of the fully phase-correlated state - coherent state $|\alpha\rangle$ is characterized by its well-defined phase and effectively accounts for all residual contributions to the detected mode, other than the dominant phase randomized fields with chaotic statistics scattered in the forward and backward directions. These include the residual scattering of the excitation laser light on atomic vapor cell windows to the detected mode, residual temporal multi-modeness due to the finite temporal resolution of the employed single-photon counting modules, and their noise - dark counts. The corresponding $g^{(2)}(0)$ can be expressed as
$g^{(2)}(0)=\frac{G^{(2)}(0)+2\bar{n}|\alpha|^{2}+|\alpha|^{4}}{\left(\bar{n}+|\alpha|^{2}\right)^{2}},$ (S7)
where $G^{(2)}(0)=2\bar{n}^{2}$ is given by Eq. (S0.2), $\bar{n}=(\bar{n}_{F}+\bar{n}_{B})$ represents the mean photon number of the scattered light and $|\alpha|^{2}$ is a mean photon number of the coherent state. Let $r=|\alpha|^{2}/\bar{n}$ denote the ratio of the mean photon numbers of phase-correlated, i.e. coherent light to the light scattered from atoms. Eq. (S7) can then be simplified to
$g^{(2)}(0)=\frac{2+2r+r^{2}}{(1+r)^{2}}.$ (S8)
Eq. (S8) accounts for the suppression of $g^{(2)}(0)$ using the single parameter - ratio $r$. For fully phase randomized contributions, $g^{(2)}(0)=2$ and $r=0$, while we obtain $g^{(2)}(0)=1$ in the limit $r\rightarrow\infty$ corresponding to observation of phase-correlated laser light.